OpenStudy (oswaldmurphy):

Given the trinomial, what is the value of the coefficient 'B' in the factored form? 2x2 − 12xy − 32y2 = 2(x − 8y)(x + By) B= ____________

OpenStudy (oswaldmurphy):

@Jadeishere

OpenStudy (mayankdevnani):

First take 2 common from the given equation

OpenStudy (oswaldmurphy):

Jade, I want you to help. You are good at explaining and dont get frustrated with me :P http://prntscr.com/c17g1k Is it 4 and 8?

OpenStudy (mayankdevnani):

Oswald nice choice mate Indeed she is a great lifesaver

OpenStudy (oswaldmurphy):

No offense D:

OpenStudy (mathstudent55):

Jade gave you a good starting point. First. factor 2 out of both sides. Then divide both sides by 2. Show what you get. Then you'll go on from there.

OpenStudy (oswaldmurphy):

Dude, you gotta understand. I am not good at this. How do i factor 2 out?

OpenStudy (mathstudent55):

On the left side, notice that every coefficient (the number that multiplies x or just a number alone) is a multiple of 2. Look below at the numbers in red. \(\color{red}{2}x^2 \color{red}{- 12}xy \color{red}{-32}y^2 = 2(x − 8y)(x + By)\) All those numbers are divisible by 2.

OpenStudy (oswaldmurphy):

Ok. I know what coefficients are. So what do we do with those numbers? Find the gcf of all of them?

OpenStudy (mathstudent55):

You divide every one of those numbers by 2. You enclose that expression in parentheses, and you write the 2 outside the parentheses. \(2(\color{red}{}x^2 \color{red}{- 6}xy \color{red}{-16}y^2 = 2(x − 8y)(x + By)\) You see how each coefficient was divided by 2, and the 2 is now outside the parentheses?

OpenStudy (mathstudent55):

Now we have a factor of 2 one each side, so we can simplify the problem by dividing the two sides by 2. \(2(x^2 - 6xy -16y^2) = 2(x − 8y)(x + By)\) Now we divide both sides by 2, and we get rid of those two factors of 2. We get this: \(x^2 - 6xy -16y^2 = (x − 8y)(x + By)\)

OpenStudy (mathstudent55):

Now we need to multiply out the right side. We have the product of two binomials on the right side, so we can use FOIL. Are you familiar with FOIL, or with multiplying binomials together?

OpenStudy (oswaldmurphy):

Remind me what Foil is please :D

OpenStudy (oswaldmurphy):

@mathstudent55 hello?

OpenStudy (jadeishere):

https://www.google.com/#q=foil+method

OpenStudy (mathstudent55):